Flamo:

The hypotenuse of a right triangle is 15 cm, and the shorter leg is 9 cm. Find the length of the other leg.

Flamo:

15^2 + 9^2 = c2

Angle:

careful!

Flamo:

?

Angle:

this time, the hypotenuse = 15 = c

Flamo:

o

Flamo:

O, I c

Flamo:

a^2 + 9^2 = 15^2?

Angle:

exactly :)

Flamo:

o

Flamo:

Where does the A go?

Angle:

we try to solve for a we want to know what a = ?

Flamo:

o

Flamo:

So its Like this then, 9^2 + 15^2 = a^2

Flamo:

To make it in a Simpler Format

Angle:

not quite a^2 + 9^2 = 15^2 subtract (9^2) from both sides so a^2 = 15^2 - 9^2

Flamo:

O

Flamo:

a^2 = 225 - 81

Flamo:

a^2 = 144

Flamo:

144 in a Square Root will be 12 Cm. :D

Angle:

yup!

Flamo:

Ye!